OpenStudy (anonymous):

Please Help me with these 2 questions I have been trying to figure them out for over 30 minutes ! 1.Use the substitution method to solve the following system of equations. 6x + 3y = -7 12x + y = 6 2.If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the second equation is substituted into the first equation. 2x - 5y = -3 6x - 3y = -3 these are my options 2(2x + 1) - 5y = -3 2(-2x - 1) - 5y = -3 2x - 5(2x + 1) = -3 2x - 5(-2x - 1) = -3 Thank you in advance!

OpenStudy (anonymous):

\[12x + y = 6 \iff y=6-12x\] is a start

myininaya (myininaya):

Take that (6-12x) and replace y in your first equation with it :)

OpenStudy (anonymous):

then \[6x + 3y = -7\] becomes \[6x+3(6-12x)=-7\]

myininaya (myininaya):

Then use the distributive property a(b+c)=ab+ac on 3(6-12x)

OpenStudy (anonymous):

to get \[6x+18-36x=-7\]

myininaya (myininaya):

Collect like terms to get...

OpenStudy (anonymous):

\[-30x+18=-7\]

myininaya (myininaya):

Now we want to isolate x...

OpenStudy (anonymous):

\[-30x=-7-18=-25\]

myininaya (myininaya):

The last thing to undo is that multiplication by -30

OpenStudy (anonymous):

\[x=-25\div -30=\frac{5}{6}\]

myininaya (myininaya):

Then you take that value you found for x and plug it into one of the equations given to find y

myininaya (myininaya):

or as satellite had said y=6-12x

myininaya (myininaya):

oThis is already solved for y all you have to do is replace x with 5/6 and simplify

OpenStudy (anonymous):

\[12x + y = 6 \] \[y=6-12(\frac{5}{6})\]

OpenStudy (anonymous):

\[y=6-2\times 5=6-10=-4\]

myininaya (myininaya):

And that is how you use the order of operations :)

myininaya (myininaya):

To simplify the value for y

OpenStudy (anonymous):

myininaya not only is the coolest person in open study, she also has the best hair do

OpenStudy (anonymous):

Thank you guys soooooooooooooooooooooooooooooooooooooooooooooooooooooooooo much ! This helped me big time and showed me exactly how to go about solving that! I really appreciate it!

OpenStudy (anonymous):

yw

myininaya (myininaya):

@satellite73 did most of the work. He is a brilliant teacher.

OpenStudy (anonymous):

gotta run, honeymooners is on no no, you did all the work, i did the donkey part (mechanics)